### v1

# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

Abstract
This work is motivated by in vivo experiments in which measurement are destructive so that the variables of interest can never be observed simultaneously when the aim is to estimate the regression coefficients of a linear regression. Assuming that the global experiment can be decomposed into sub experiments (corresponding for example to different doses) with distinct first moments, we propose different estimators of the linear regression which take account of that additional information. We consider estimators based on moments as well as estimators based optimal transport theory. These estimators are proved to be consistent as well as asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some particular cases, and specific bootstrap approaches are developed to build confidence intervals for the estimated parameter. A Monte Carlo study is conducted to assess and compare the finite sample performances of the different approaches.
1 Motivation and context
In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted. In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect from different perspectives. In particular, such variables of interest often include a macroscopic biomarker only available through in vivo data on the one hand, and a microscopic biomarker that can also be observed on a cellular level. The interest in this case lies in predicting the former with the latter. For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest may manifest in the form of lesions, that are quantified in a form a certain macroscopic biomarker, and the levels of a some predictor such as gene expression. These measures often require sacrificing the animal. As a result, the quantities of interest cannot be observed on the same animals. Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component (experimental condition). An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
1.1 Example of an in vivo experiment
Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1. In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearing of radio-induced adverse effects. The latter are assessed by measuring septal thickening, a macroscopic biomarker of radio-induced adverse effects in the lungs. The other variable that is measured with the purpose of predicting the adverse effect related variable is the expression of multiple pro-inflammatory genes. As shown in Figure 1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness. This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 2, one may suspect a correlation or even a linear relationship between the variables. In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem. In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken. Thus, the categorical variable indication the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
1.2 Existing research
The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here. As a result, any approaches applying related frameworks such as multiple imputation in the context of data fusion are unsuitable for our application. For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest. This corresponds to the properties of the in vivo data described in Section 1.1. However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The groups variable available through in vivo experiments may present in a continuous form, but the presence of such categories as control and sham makes it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
2 Estimating the linear relation between variables that are never jointly observed
2.1 Problem and presentation of the different identification approaches
We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$ and suppose that the following linear regression hold
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
[…]
In the absence of additional information and without any strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified. Indeed, if for example $X_{1}$ is centered with symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to sign change since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
To deal with this identification issue, we consider that we can perform different experiments in which the mean value of $X$ is allowed to vary. For that, we suppose that there are $K$ groups (corresponding to $K$ different experiments), defined by a discrete variable $G$ taking values in $\{1,\cdots,K\}$ observed simultaneously with $Y$ and with $X$. This means that we have now access to $(\mathbf{X},G)$ and $(Y,G)$ but not to $(\mathbf{X},Y,G)$. We also suppose that $\epsilon$ is independent of $G$.
Given $G=k$, for $k=1,\ldots,K$, we denote by $\mu_{Y}^{k}={\mathbb{E}}(Y|G=k)$ and $\mu_{X_{j}}^{k}={\mathbb{E}}(X_{j}|G=k)$, $j=1,\ldots,d$, the expected values within each group.
We now present two different approaches developed to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking account of the additional information given by the discrete variable $G$.
2.1.1 Moment approach
A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
since the residual term $\epsilon$ is supposed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
We introduce the following assumption, guaranteeing the identifiability of the model parameters:
$\mathbf{H}_{1}$ $\quad\mbox{rank}(\boldsymbol{\mu}_{1,X})=d+1,$
meaning that there are at least $K\geq d+1$ groups and that the $d+1$ column vectors of $\boldsymbol{\mu}_{1,X}$ span a vector space of dimension $d+1$ in $\mathbb{R}^{K}$.
Lemma 2.1.
If model (1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
$\boldsymbol{\beta}$ $=\left(\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{Y}.$
Additionally, the noise variance $\sigma_{\epsilon}^{2}$ satisfies
$\sigma_{\epsilon}^{2}$ $=\sigma_{Y}^{2}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}\boldsymbol{\beta}_{-0},$
where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
The proof of Lemma 2.1 is direct and thus omitted.
2.1.2 Optimal transport and minimum Wasserstein distance approach
The second approach is based on optimal transport, in particular on the idea of estimating the linear transformation of the distribution of $\mathbf{X}$ that is the closest to that of $Y$ with respect to the Wasserstein distance (see Panaretos and Zemel, (2019) for a general introduction for statisticians).
The optimal transport map T between Gaussian measures on $\mathbb{R}^{d}$ is linear, and the Wasserstein distance of order 2 between two Gaussian distributions $D_{1}$ and $D_{2}$, with $D_{1}\sim\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Gamma}_{1})$ and $D_{2}\sim\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Gamma}_{2})$, is equal to
$W_{2}^{2}(D_{1},D_{2})$ $=\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+\mbox{tr}\left(\boldsymbol{\Gamma}_{1}+\boldsymbol{\Gamma}_{2}-2\left(\boldsymbol{\Gamma}_{2}^{1/2}\boldsymbol{\Gamma}_{1}\boldsymbol{\Gamma}_{2}^{1/2}\right)^{1/2}\right),$
[…]
If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
$W_{2}^{2}(D_{\gamma},D_{Y})$ $=\mathbb{E}\left[W_{2}^{2}(D_{\gamma},D_{Y})|G\right]$
$=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\alpha_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\epsilon}^{2}}\right)^{2}\right],$
where $\pi_{k}=\mathbb{P}[G=k]$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$.
We introduce the following loss criterion
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by a Gaussian noise, with variance $\sigma^{2}$. We state, without proof, the following Lemma ensuring that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions.
Lemma 2.2.
If model (1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
2.2 Sampled data and estimators
We suppose that experiments are made for $K\geq 2$ different groups and that for each group $k$, for $k=1,\ldots,K$, we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$. For each unit $i=1,\ldots,n_{x}^{k}$ from group $k$, the vector of covariates is denoted by $\mathbf{X}_{i}^{k}=(X_{1,i},\ldots,X_{d,i})$. We also define $N_{x}=\sum_{k=1}^{K}n_{x}^{k}$ and $N_{y}=\sum_{k=1}^{K}n_{y}^{k}$, the total number of observations of the response $Y$ and the covariates $X_{1},\ldots,X_{d}$.
As shown in Lemma 2.1 and Lemma 2.2, identification of parameter $\boldsymbol{\beta}$ relies on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$. For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical means within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$. We also define the overall empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$.
We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.
2.2.1 Moment estimators
If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of identification equations given in Lemma 2.1,
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$
where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$. We consider in the following a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$. The weighted moment estimator $\boldsymbol{\beta}$ is defined as the minimizer of
$\psi(\boldsymbol{\gamma})$ $=\sum_{k=1}^{K}w_{k}\left[\widehat{\mu}_{Y}^{k}-\left(\gamma_{0}+\sum_{j=1}^{d}\gamma_{j}\widehat{\mu}_{X_{j}}^{k}\right)\right]^{2},$
which is unique if $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank and defined by
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered just before corresponds to equal weights $w_{k}=K^{-1}$.
We can then define the following estimator of the noise variance,
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (5)
2.2.2 Minimum Wasserstein distance estimators
Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (6)
Note that in absence of an a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
2.3 Consistency and asymptotic distribution
To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in previous Section, we suppose that the number $K$ of groups is kept fixed and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity. This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$ the smallest sample size among all experiments is also supposed to tend to infinity.
Under general moment conditions, without any particular assumption on the probability law of $Y$ and $\mathbf{X}$ we get in the following two Lemmas that the moment estimators and the Wasserstein distance based estimators are consistent. The proofs are postponed to the Appendix.
Lemma 2.3.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (4) and (5) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
For the Wasserstein minimum distance estimator, since there is no explicit expression for the estimators, a compactness assumption is also made to obtain the consistency.
Lemma 2.4.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (6) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
As far as the asymptotic distribution of the estimators is concerned, and for sake of simplicity and lighter notations, we suppose from now on that the number of experiments is the same for all groups and all variables, that is to say $n=n_{y}^{1}=\cdots n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
Proposition 2.1.
Assume that the assumptions of Lemma 2.3 are fulfilled. Then as $n$ tends to infinity,
$\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right)$
where the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$ is given in the proof.
This result is based on the CLT for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making obtaining the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ a difficult task when $d>1$.
Remark 2.1.
The weak convergence toward a Gaussian distribution presented in Proposition 2.1 would remain true, at the expense of heavier notations and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (7)
and $\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})\to\infty$.
The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in the Appendix (see Theorem A.4). Note that since we estimate simultaneously $\beta$ and $\sigma_{\epsilon}^{2}$, an additional condition on the existence of the moments of order four for the covariates is required.
Proposition 2.2.
If model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
$\sqrt{n}\left(\begin{pmatrix}\widehat{\boldsymbol{\beta}}^{W}\\ \widehat{\sigma}^{2,W}_{\epsilon}\end{pmatrix}-\begin{pmatrix}\boldsymbol{\beta}\\ \sigma^{2}_{\epsilon}\end{pmatrix}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{W}\right),$
for some covariance matrix $\boldsymbol{\Gamma}_{W}$.
Similarly to $\boldsymbol{\Gamma}_{\beta_{M}}$, the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{W}$ is almost impossible to explicitly derive manually, with the exception of some particularly simple cases.
3 Bootstrapping for confidence intervals
Since, as noted in the previous section, it is complicated to compute explicitly the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$.
Our bootstrap procedure takes account of the independence between the different groups $k=1,\ldots,K$, as well as the independence of inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and output $Y^{k}$ within each group, meaning more formally that, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (4) and (6).
To build confidence sets for the components of $\boldsymbol{\beta}$ based on previous bootstrap procedure, the bootstrap percentile technique, described in Chapter 4 of Shao and Tu, (1995), is simple to use.
It can be noted that our estimators are smooth functions of sample means so that classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3). For simplicity, as in Proposition 2.1, we suppose that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$. Because of the considered experimental design, our global "empirical distribution" is made of products of marginal empirical distributions, the bootstrap for means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work. Then, the application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that bootstrap percentile method gives consistent confidence sets for each component of $\boldsymbol{\beta}$.
Proposition 3.1.
Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and hypothesis $\mathbf{H}_{1}$ is fulfilled. Then as $n\to+\infty$, the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and compactness assumption.
Proposition 3.2.
Suppose that the assumptions of Proposition 2.2 as fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
Propositions 3.1 and 3.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
4 A simulation study
Some simulations are performed in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
4.1 Simulation design
4.1.1 General setting
Let $K\in\mathbb{N}$ be the number of subpopulations, $g_{k}=\{0,1,\dots,k-1\}$ be the level of the grouping variable for subpopulation $k\in\{1,\dots,K\}$, and the number of animals observed per group was choose for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$. For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$ we choose the predictor variable $X^{k}_{i}$ in the Gaussian univariate setting:
$X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X,k}^{2})\text{, where }\mu_{X}^{k}=\Delta_{\mu}g_{k}+C_{\mu},$
where $C_{\mu}\in\mathbb{R}$ is the global mean scale parameter, and $\Delta_{\mu}\in\mathbb{R}$ is the parameter that determines the difference between groups in terms of the mean. We simulate the predicted variable independently as follows, with regression parameters $\beta_{0}$ and $\beta_{1}$:
$Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}\text{, where }X^{\prime k}_{i}\sim X^{k}_{i}\text{ and }\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon,k}^{2}).$
It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}\in\mathbb{N}$ times. For each simulation, $N_{boot}\in\mathbb{N}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated. Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e. the proportion of intervals not including 0. Along with the bootstrap procedure, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
$\frac{\widehat{\beta}_{j}-\beta_{j}}{SE(\widehat{\beta_{j}})}\rightsquigarrow t_{K-(d+1)}\text{ for }j\in\{0,\dots,d\}.$
4.1.2 Parameter levels
We fix the following parameters throughout all simulations: the regression parameters $\beta_{0}=1$ and $\beta_{1}=2$, the number of simulations $N_{sim}=500$, the number of bootstrap samples $N_{boot}=500$, the location parameters of the predictor variable $\Delta_{\mu}=1$ and $C_{\mu}=10$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired in case of a small number of animals, which is often the case for real experimental data. We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution. Additionally, the parameter $\sigma^{2}_{X,k}$ can be adjusted to control the extent to which the observations per group can be easily distinguish one from another. We set $\sigma_{X,1}^{2}=\dots=\sigma_{X,K}^{2}$, and $\sigma^{2}_{X,k}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e. $\rho=\cfrac{\sigma_{Y,k}}{\sigma_{\epsilon,k}}$. The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon,k}^{2}$ is motivated by the fact that $\sigma_{Y,k}$ depends on $\sigma_{X,k}$, hence the same level of $\sigma_{\epsilon,k}$ cannot be interpreted the same way for different values of $\sigma_{X,k}$. The variance of the noise can be expressed as follows: $\sigma_{\epsilon,k}^{2}=\cfrac{\beta_{1}^{2}\sigma_{X,k}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely the very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 3.
Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X,1}^{2}=\dots=\sigma_{X,4}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
4.2 Results
The results of the simulation study are presented in Figure 4, namely the obtained coverage rates for the linear regression slope confidence intervals on Figure 3(a), the amplitudes on Figure 3(b), and the associated test powers on Figure 3(c). Every figure contains four tables, each corresponding to a combination of parameters with respect to the number of animals $N_{anim}$ and the number of groups $K$. The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X,k}$ and the noise to signal ratio $\rho$. Finally, the lines of each table indicate the method used to estimate the confidence intervals: the proposed method of moments and optimal transport based bootstrap estimators, and the classical linear regression procedure based on estimated means within each group.
Figure 4: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination. The method labels on the left: "mm (boot)" stands for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
In general, it can be observed that the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, but with lower coverage rate. Whereas the coverage rates in all cases remain within 91-96%, the extent to which the amplitudes are smaller and the powers are bigger is significantly more important for the bootstrap estimators in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.This trend of further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches. Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present. Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X,k}$ and higher $\rho$. In most case, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high dispersion and high noise level, with few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes. On the other hand, we can also distinguish two case where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals. This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X,k}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
5 Application to real data
In order to illustrate the proposed estimators on real data, we studied the data from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume. In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy. For one cohort, the expressions of three pro-inflammatory genes (IL1 $\alpha$ and $\beta$, IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a macroscopic biomarker of radio-induced pulmonary lesions. In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung). The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as predictors and the septal thickening as as outcome. Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
Method of moments Optimal transport Lin. Reg. on means
(bootstrap) (bootstrap) (Student)
Loc. Vol. Gene ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif.
IL1b -0.35 (-1.03, 0.4) -0.19 (-0.44, 0.23) -0.35 (-1.99, 1.28)
1 mm IL6 0.43 (-0.23, 1.2) 0.2 (-0.2, 0.84) 0.43 (-1.32, 2.17)
Ipsilateral TNF 0.2 (-0.23, 0.84) 0.16 (-0.18, 0.84) 0.2 (-1.26, 1.66)
lung IL1b 0.9 (0.0, 2.11) 1.05 (0.01, 1.44) 0.9 (-0.43, 2.24)
3 mm IL6 0.05 (-0.34, 0.46) 0.06 (-0.33, 0.49) 0.05 (-1.65, 1.76)
TNF 0.65 (-0.12, 1.6) 0.63 (-0.12, 1.46) 0.65 (-1.57, 2.87)
IL1b 2.31 (-1.73, 3.87) 0.82 (-0.48, 0.94) 2.31 (-0.73, 5.35)
1 mm IL6 0.85 (-0.7, 2.38) 0.61 (-0.56, 1.6) 0.85 (-3.75, 5.45)
Irradiated TNF 0.85 (-0.6, 2.3) 0.69 (-0.54, 1.53) 0.85 (-3.96, 5.66)
patch IL1b 2.85 (0.84, 5.14) 2.59 (0.73, 3.58) 2.85 (-1.07, 6.78)
3 mm IL6 1.35 (0.22, 2.47) 1.3 (0.21, 2.26) 1.35 (-1.8, 4.5)
TNF 3.81 (1.01, 6.37) 3.37 (0.99, 5.33) 3.81 (-1.86, 9.48)
IL1b -0.92 (-2.2, 1.06) -0.42 (-0.89, 0.51) -0.92 (-3.97, 2.14)
0 mm IL6 2.23 (-1.99, 2.28) 0.83 (-0.83, 0.97) 2.23 (-2.65, 7.12)
TNF 2 (-2.0, 2.82) 0.88 (-1.0, 1.11) 2 (-1.72, 5.72)
IL1b 0.63 (-0.82, 2.32) 0.23 (-0.33, 0.56) 0.63 (-2.95, 4.21)
1 mm IL6 1.03 (-0.57, 2.19) 0.46 (-0.3, 1.0) 1.03 (-0.88, 2.94)
Right TNF 1.05 (-0.47, 2.43) 0.66 (-0.31, 1.26) 1.05 (-1.01, 3.11)
lung IL1b 1.07 (0.17, 1.97) 0.92 (0.17, 1.13) 1.07 (-1.96, 4.1)
3 mm IL6 0.3 (-1.45, 1.12) 0.37 (-0.74, 0.86) 0.3 (-4.94, 5.53)
TNF 2.02 (0.27, 4.1) 1.05 (0.04, 1.44) 2.02 (0.03, 4.02)
IL1b 6.41 (-11.6, 15.4) 1.62 (-2.28, 3.1) 6.41 (-16.7, 29.52)
7 mm IL6 -0.58 (-1.26, 0.09) -0.6 (-1.23, 0.1) -0.58 (-6.23, 5.08)
TNF 0.51 (-1.7, 3.86) 0.43 (-1.45, 1.89) 0.51 (-23.21, 24.22)
Whole IL1a 2.93 (-1.18, 6.63) 2.47 (-1.49, 3.02) 2.93 (-11.08, 16.94)
thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
irradiation TNF 2.35 (0.57, 4.73) 1.67 (0.53, 1.88) 2.35 (-3.86, 8.57)
The linear regression parameters were estimated with three estimators the same way it was done in the simulation study in Section 4.2. The results are presented in Table 1. The focus is placed on estimating the slope parameter $\beta_{1}$ in particular. The table contains the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
Figure 5: Distribution of the data, collected from the irradiated patch under SBRT with 7 mm beam size: the expression of the gene TNF on the left, and septal thickness on the right. The measurements were made 30, 45 and 60 days after irradiation.
Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance. On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (two out of three genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result. On the other hand, we also expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch. This is only the case for the bootstrap estimators, but not for the naive one. This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
For the irradiated patch, as well as in all other cases with identified significance, it is only the case for the beam size of 3 mm, the results that is consistent with literature, indicating it as the beam size starting with which the long-term lesions start appearing (Bertho et al.,, 2020). Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm. However, none of the genes has been identified as significantly linked to lesions in the case of 7 mm. This may be caused by the relationship being of non-linear nature, as appears to be the case for the gene TNF, presented in Figure 5.
Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend. These values are in general bigger in case of the whole-thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, which is also in line with biological knowledge, since the genes are expected to be more strongly up-regulated in the areas of injury than further away. This effect is illustrated in Figure 6 on the example of the linear model prediction made for the gene IL6 .
Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
6 Discussion
This work focuses on a statistical framework designed for extracting dependencies form in vivo experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed. In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility. Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly. The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied. For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects. To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model. The optimal transport estimator appears promising in this context given the fact that Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions. Concerning a likelihood-based approach, the joint likelihood to be maximized can be written in a factorized form due to the independence of $X$ and $Y$: $L(x,y;\beta_{0},\beta_{1},\mu_{x}^{k},\sigma_{\epsilon},\sigma_{x})=L_{X}(x;\mu_{x},\sigma_{x})L_{Y}(y;\beta_{0},\beta_{1},\mu_{x},\sigma_{\epsilon},\sigma_{x})$ for a group $k$ in the univariate setting. The marginal likelihoods can be expressed explicitly, assuming that the model (1) is true, and that $X|(G=k)\sim\mathcal{N}(\mu_{x}^{k},\sigma_{x,k}^{2})$, and thus $Y|(G=k)\sim\mathcal{N}(\beta_{0}+\beta_{1}\mu_{x}^{k},\sigma_{\epsilon}^{2}+\beta_{1}^{2}\sigma_{x,k}^{2})$, with the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$.
Finally, it would be of interest to work on improving theoretical properties of our estimators, namely correcting the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023). Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
Funding
This work is supported by the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and by Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
Conflict of Interest: none declared.
Acknowledgements: The authors warmly thank Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
Appendix A Some classical theorems in asymptotic statistics
A proof of the classical continuous mapping theorem can be found in van der Vaart, (1998) (Theorem 2.3).
Theorem A.1.
(Continuous mapping theorem).
Let $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ be continuous at every point of $C$ such that $\mathbb{P}[X\in C]=1$.
If the sequence of random variables $(X_{n})_{n\geq 1}$ converges in distribution (resp. probability, resp. almost surely) to $X$ then $(g(X_{n}))_{n\geq 1}$ converges in distribution (resp. probability, resp. almost surely) to $g(X)$.
We also recall some well known results that are useful to show the consistency of estimators $\widehat{\theta}_{n}$ defined as the minimizers of functionals $Q_{n}(\theta)$ which have some regularity properties at the limit.
Theorem A.2.
(Lemma 2.9 in Newey and McFadden, (1994))
Suppose that $\theta\in\Theta$ and $\Theta$ is compact, $Q_{0}(\theta)$ is continuous and $\forall\theta\in\Theta$, $Q_{n}(\theta)\to Q_{0}(\theta)$ in probability as $n$ tends to infinity. If there is $\alpha>0$ and $B_{n}=O_{p}(1)$ such that
$\forall(\widetilde{\theta},\theta)\in\Theta\times\Theta,\ |Q_{n}(\widetilde{\theta})-Q_{n}(\theta)|\leq B_{n}\|\widetilde{\theta}-\theta\|^{\alpha}$
then
$\sup_{\theta\in\Theta}\left|Q_{n}(\theta)-Q_{0}(\theta)\right|\to 0\ \mbox{ inprobability.}$.
Theorem A.3.
(Theorem 2.1 in Newey and McFadden, (1994))
Suppose that $\theta\in\Theta$ and $\Theta$ is compact, $Q_{0}(\theta)$ is continuous $\forall\theta\in\Theta$. If $Q_{0}(\theta)$ is uniquely maximized at $\theta_{0}$ and, as $n$ tends to infinity, $\sup_{\theta\in\Theta}\left|Q_{n}(\theta)-Q_{0}(\theta)\right|\to 0$ in probability, then $\widehat{\theta}_{n}\to\theta_{0}$ in probability.
Under additional hypotheses, we also get the asymptotic normality of the sequence of estimators $\widehat{\theta}_{n}$ of $\theta_{0}$. we denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of functional $Q_{n}$ evaluated at $\theta$.
Theorem A.4.
(Theorem 3.1 in Newey and McFadden, (1994))
Suppose that $\widehat{\theta}_{n}\to\theta_{0}$ in probability, (i) $\theta_{0}$ is an interior point of $\Theta$, (ii) $Q_{n}(\theta)$ is twice differentiable in a neighborhood $\mathcal{N}$ of $\theta_{0}$, (iii) $\sqrt{n}\nabla_{0}Q_{n}(\theta_{0})\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Sigma}\right)$, (iv) there is $\mathbf{H}(\theta)$ continuous at $\theta_{0}$ and $\sup_{\theta\in\mathcal{N}}\|\nabla_{00}Q_{n}(\theta)-\mathbf{H}(\theta)\|\to 0$ in probability (v) $\mathbf{H}=\mathbf{H}(\theta_{0})$ is non singular. Then
$\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\rightsquigarrow\mathcal{N}\left(0,\mathbf{H}^{-1}\boldsymbol{\Sigma}\mathbf{H}^{-1}\right)$
We also recall the central limit theorem for bootstrap means (see Theorem 23.4 in van der Vaart, (1998) for a proof).
Theorem A.5.
(CLT for bootstrap means)
Let $X_{1},X_{2},\ldots$ be i.i.d. random vectors with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Gamma}$. Then conditionally on $X_{1},X_{2},\ldots$, for almost every sequence $X_{1},X_{2},\ldots$
$\sqrt{n}\left(\overline{X}^{*}_{n}-\overline{X}_{n}\right)\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}\right)$
where $\overline{X}_{n}$ is the empirical mean and $\overline{X}^{*}_{n}$ is the empirical mean of $n$ independent observations drawn from the empirical distribution.
Appendix B Proofs
Proof.
of Lemma 2.3
First note that the assumptions ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$ ensure the existence of $\sigma_{Y}^{2}$ and $\boldsymbol{\Gamma}_{X}$. From the law of large numbers, we have that for all $k\in\{1,\ldots,K\},\ \widehat{\boldsymbol{\mu}}_{1,X}^{k}\to\boldsymbol{\mu}_{X}^{k}$ and $\widehat{\mu}_{Y}^{k}\to\mu_{Y}^{k}$ in probability when $n_{\min}$ tends to infinity. We deduce from the continuous mapping theorem that $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ and $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}$ in probability. Under hypothesis $\mathbf{H}_{1}$, the inverse being continuous in a neighborhood of $\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ another application of the continuous mapping theorem gives that $\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}$ and
$\widehat{\boldsymbol{\beta}}^{M}=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}=\boldsymbol{\beta}$
in probability as $n_{\min}$ tends to infinity. The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. ∎
Proof.
of Lemma 2.4
The proof is based on Lemma 2.9 and Theorem 2.1 in Newey and McFadden, (1994), which are recalled in Appendix A. The law of large numbers and the continuous mapping theorem give us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity. Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$. We have,
$\left|(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}-(\widehat{\mu}_{Y}^{k}-\alpha_{0}-\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}\right|$ $=\left|(\boldsymbol{\alpha}-\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\left(2\widehat{\mu}_{Y}^{k}-(\boldsymbol{\alpha}+\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\right)\right|$
$\leq\left\lVert\boldsymbol{\alpha}-\boldsymbol{\gamma}\right\rVert A_{n}^{k},$
with Cauchy-Schwarz inequality and $A_{n,k}=\mathcal{O}_{p}(1)$ because $\|\widehat{\boldsymbol{\mu}}_{X}^{k}\|=O_{p}(1)$, $\widehat{\mu}_{Y}^{k}=O_{p}(1)$ and for some constant $C_{1}$ that does not depend on $\boldsymbol{\alpha}$ and $\boldsymbol{\gamma}$, $\|\boldsymbol{\alpha}+\boldsymbol{\gamma}\|\leq C_{1}<\infty$ because $\Theta$ is supposed to be compact. On the other hand, we have
$\left|\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right)^{2}-\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}\right)^{2}\right|$
$=\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|\left(2\widehat{\sigma}_{Y,k}+\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}+\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right)$
$=\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|O_{p}(1)$
since $\Theta$ is compact and $\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\|_{sp}=O_{p}(1)$, where .$\|.\|_{sp}$ denotes the spectral norm. Because $\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}=\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)+\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}$ we have, for some constant $C_{2,k}>0$,
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (8)
Using now the fact that function $x\mapsto\sqrt{x}$ is concave and differentiable, we have for $x>0$ and $y>0$ that $\sqrt{y}\leq\sqrt{x}+\frac{y-x}{2\sqrt{x}}$. Thus, if $y>x>0$ then $0<\sqrt{y}-\sqrt{x}\leq\frac{y-x}{2\sqrt{x}}$ and if $x>y>0$, then $0<\sqrt{x}-\sqrt{y}\leq\frac{x-y}{2\sqrt{y}}$. Consequently, we have $|\sqrt{y}-\sqrt{x}|\leq\frac{|x-y|}{2\min(\sqrt{x},\sqrt{y})}$ and we deduce that,
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (9)
where $B_{n}^{k}=O_{p}(1)$. Combining previous inequalities, we get
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (10)
with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right)=O_{p}(1)$. As a result, it can be deduced from Lemma 2.9 in Newey and McFadden, (1994) that
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
We conclude the proof by recalling that $\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ attains its unique minimum at $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ if assumption $\mathbf{H}_{1}$ is fulfilled, so that $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})\to(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). ∎
Proof.
of Proposition 2.1
The central limit theorem applies directly to the independent sequences of independent random variables $(\mathbf{X}_{1}^{1},\cdots,\mathbf{X}_{n}^{1}),\ldots,(\mathbf{X}_{1}^{K},\cdots,\mathbf{X}_{n}^{K})$ and $(Y_{1}^{1},\cdots,Y_{n}^{1}),,\ldots,(Y_{1}^{K},\cdots,Y_{n}^{K})$ so that, as $n$ tends to infinity
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (17)
where $\boldsymbol{\Gamma}_{\mu}$ is a block diagonal matrix, with diagonal elements $(\boldsymbol{\Gamma}_{X}^{1},\ldots,\boldsymbol{\Gamma}_{X}^{K},\sigma_{Y,1}^{2},\ldots,\sigma_{Y,K}^{2})$, with $\boldsymbol{\Gamma}_{X}^{k}=\mbox{Var}(\mathbf{X}|G=k)={\mathbb{E}}\left(\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}\right)-\boldsymbol{\mu}_{X}^{k}(\boldsymbol{\mu}_{X}^{k})^{\top}$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$. Consider the application $g:\mathbb{R}^{dK+K}\to\mathbb{R}^{d+1}$ defined by
$g(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})=\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}.$
Application $g$ is differentiable at $\boldsymbol{\theta}=(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, with non null Jacobian matrix denoted by $\mathbf{J}_{\theta}$ (see Chapter 8 and more particularly Theorem 8.3 in Magnus and Neudecker, (2019)). The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) permits to get the asymptotic normality convergence result,
$\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right),$
where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. ∎
Proof.
of Proposition 2.2
The proof consists in checking the different points of Theorem A.4. Point (i) is satisfied by the hypotheses, and the point (ii) follows directly from the fact $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$. To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$,
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (21)
Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (17) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled. To prove that (iv) also holds, consider the Hessian matrix of functional $\varphi_{n}$, evaluated at $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$:
$\nabla_{00}\varphi_{n}$ $=\begin{pmatrix}2&2\left(\sum_{k=1}^{K}\pi_{k}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)^{\top}&0\\ 2\sum_{k=1}^{K}\pi_{k}\widehat{\boldsymbol{\mu}}_{X}^{k}&\widehat{\mathbf{H}}(\boldsymbol{\beta}_{-0})&\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\\ 0&\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}&\frac{1}{2}\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\end{pmatrix},$
where
$\widehat{\mathbf{H}}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\bigg{[}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\bigg{[}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\bigg{]}$
$+\widehat{\boldsymbol{\mu}}_{X}^{k}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}\right)^{\top}+\widehat{\boldsymbol{\Gamma}}_{X}^{k}\bigg{]}.$
By similar arguments as those used to show that $\varphi_{n}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ converges in probability to $\varphi(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we deduce that $\nabla_{00}\varphi_{n}$ converges in probability to some matrix $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, defined as follows
$\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ $=\begin{pmatrix}2&2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}&0\\ 2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\mathbf{H}(\boldsymbol{\beta}_{-0})&\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\\ 0&\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}&\frac{1}{2}\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\end{pmatrix}$
where
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\bigg{(}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\bigg{[}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)\boldsymbol{\Gamma}_{X}^{k}\bigg{]}$
$+\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\boldsymbol{\Gamma}_{X}^{k}\bigg{)}.$
We now must check that $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is a positive definite matrix. For that we show that at the minimizer value $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ its determinant is strictly positive. We first note that $\sigma_{Y,k}=\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{1/2}$ so that $\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}=\frac{1}{\sigma_{Y,k}^{2}}$ and $\mathbf{H}(\boldsymbol{\beta}_{-0})$ can be written in a simpler form,
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (22)
which is a definite positive matrix under hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (28)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (28) is strictly positive. We now have to show that
$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$ (31)
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (32)
is a positive matrix. We can remark that by Cauchy Schwarz inequality, for $\mathbf{u}\in\mathbb{R}^{d}$,
$\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}\mathbf{u}=\left(\sum_{k=1}^{K}\pi_{k}\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{k}\right)^{2}$ $\leq\sum_{k=1}^{K}\pi_{k}\left(\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{k}\right)^{2}$
$=\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}\right)\mathbf{u}$
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (32) is handled the same way. We have
$\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\mathbf{u}$ $=\left(\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\right)^{2}$
$\leq\sum_{k=1}^{K}\sqrt{\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}^{2}\sum_{k=1}^{K}\left(\sqrt{\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}^{2}\mathbf{u}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{2}$
$=\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\mathbf{u}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\mathbf{u},$
and consequently the determinant of $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is strictly positive. To finish the proof, it remains to check that in a neighborhood $\mathcal{N}$ of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we have
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}}\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0\ \mbox{ in probability.}$.
This is a direct consequence of the continuous mapping theorem, which gives us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}$, $\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0$ in probability, and the fact that third order partial derivatives of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ are bounded in probability for $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ so that Theorem A.2 can apply. ∎
Proof.
of Proposition 3.1
The fact that the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ is a direct consequence of Theorem 3.1 in Shao and Tu, (1995), noting that
$\widehat{\boldsymbol{\beta}}^{M}=g(\widehat{\boldsymbol{\mu}}_{X}^{1},\ldots,\widehat{\boldsymbol{\mu}}_{X}^{K},\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$
is a continuously differentiable function of means at $(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The fact that confidence sets based on the percentile approach are consistent is proved by checking the assumptions in Theorem 4.1 (iii) Shao and Tu, (1995), namely the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is consistent, $\widehat{\boldsymbol{\beta}}^{M}$ is consistent (Lemma 2.3), with asymptotic Gaussian distribution (Proposition 2.1). ∎
Proof.
of Proposition 3.2
We denote by $\boldsymbol{\theta}_{0}=(\boldsymbol{\beta},\sigma_{\epsilon}^{2,W})$ the vector of true parameters, by $\widehat{\boldsymbol{\theta}}=(\boldsymbol{\beta}^{W},\sigma_{\epsilon}^{2})$ the sequence of minimum Wasserstein distance estimators and by $\boldsymbol{\theta}^{*}=(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ bootstrap estimators of $\boldsymbol{\theta}_{0}$. The vector of parameters $\boldsymbol{\theta}^{*}$ is the minimizer of functional $\varphi_{n}^{*}$ defined as follows,
$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (33)
We first show with arguments similar to those employed in the proof of Lemma 2.4, that $\boldsymbol{\theta}^{*}$ is a consistent estimator for $\boldsymbol{\theta}_{0}$, based on the fact that $\varphi_{n}^{*}$ is a smooth function converging to $\varphi$ and the sample mean theorem for bootstrap (see for example Theorem 23.4 in van der Vaart, (1998)). Indeed, we first recall that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity and
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (34)
Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$. We also have, as in (10), where empirical means are replaced by bootstrap means,
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (35)
for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$. As a result, we deduce from Lemma 2.4, inequality (34) and Lemma 2.9 in Newey and McFadden, (1994) that
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
We conclude that $\boldsymbol{\theta}^{*}\to\boldsymbol{\theta}_{0}$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). We now prove that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have the same asymptotic distribution. By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (36)
where $\overline{\boldsymbol{\theta}}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\widehat{\boldsymbol{\theta}}$. We have a similar expansion for boostrap estimators, as well as
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (37)
where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (36) and (37), we deduce
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$
$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (38)
Noticing that $\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})$ and $\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})$ both tend in probability to the same limit $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ and we have, with similar arguments as those used in the proof of Proposition 2.2, that $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})$ is $O_{p}(n^{-1/2})$, it can be deduce that
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (39)
Using arguments similar to those employed in the expansion of $\nabla\varphi_{n}$ in the proof of Proposition 2.2, we make appear the difference between bootstrap means and empirical means or a differentiable functional of these quantities:
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (45)
which satisfies the central limit theorem for bootstrap means, or the Delta method for bootstrap estimators (see the Appendix as well as Theorem 23.4 and Theorem 23.5 in van der Vaart, (1998)). Consequently, $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$ and $\nabla\varphi_{n}(\boldsymbol{\theta}_{0})-\nabla\varphi(\boldsymbol{\theta}_{0})$ have the same asymptotic distribution. By Slustky’s theorem, the asymptotic distribution of $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ is thus the same as the asymptotic distribution of $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\sqrt{n}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$, and we can conclude that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have also the same asymptotic Gaussian distribution. ∎